Further to last nights scheduled maintenance at Blythe Valley Park we have experienced some teething problems today that has affected most tenants at that site. These issues boiled down to: 1) A duplex mismatch on one of our backhaul circuits into the site. 2) A number of tenants using out of date and deprecated nameservers.… Read More »
From midnight this evening LINX will be migrating our LINX Brocade connection to the new Juniper network. There maybe some small disruption to service on routes we recieve via LINX Brocade.
We are receiving reports of users not being able to access certain external websites this morning, such as The Guardian, Daily Mail, these websites predominently appear to be hosted with Akamai. We are investigating the problems and will report back shortly. UPDATE: We have narrowed this down to an Akamai issue on their LONAP port,… Read More »
24th October 2011 – 19:00 to 00:00 During the above time slot we will be carrying out the following maintenance at Blythe Valley Park POP: 1) Scheduled replacement of UPS batterys and monitoring equipment. 2) Replacement of core switching and routing infrastructure. This upgrade will replace aging equipment and give us the ability to deliver… Read More »
8/10/2011 -13:00 to 15:00 One of our switches at Netplan has developed a problem. It is currently carrying traffic without any problems however we are concerned it will deteriorate. During the above slot we will be checking and more than likely rebooting the switchin possibly leading to some packet loss while the network reconverges. We… Read More »
4th October 2011 – 22:00 to 01:00 During the above timeslot we will be carrying out further configuration changes at our London core, these changes are part of our program of network improvements. Interuption to service should be minimal, and the slot is ‘at-risk’.